# EMBRACE: Shaping Inclusive Opinion Representation by Aligning Implicit Conversations with Social Norms

### v1

@@ C.1 Positive Unlabeled Online Learning
The core implementation is derived from the fairness setting proposed by (Jung and Wang, 2024b). We modified the data preprocessing to utilize SBERT. Also, we alter the group’s definition to be represented as Implicit and Explicit.
@@
Equal Opportunity is a group fairness criterion that requires models to equalize the true positive rate (TPR) across groups defined by a sensitive attribute (Jung and Wang, 2024b) in our study we redefine that to be linked with (explicit and implicit opinion expression). Mainly, this constraint ensures that among examples who truly belong to the positive class (), the probability of being correctly classified as positive is the same across implicit and explicit groups ( and ):

### v2

@@ C.1 Positive Unlabeled Online Learning
The core implementation is derived from the fairness setting proposed by (Jung2024-pb). We modified the data preprocessing to utilize SBERT. Also, we alter the group’s definition to be represented as Implicit and Explicit.
@@
Equal Opportunity is a group fairness criterion that requires models to equalize the true positive rate (TPR) across groups defined by a sensitive attribute (Jung2024-pb) in our study we redefine that to be linked with (explicit and implicit opinion expression). Mainly, this constraint ensures that among examples who truly belong to the positive class (), the probability of being correctly classified as positive is the same across implicit and explicit groups ( and ):